563blackghost (563blackghost):

Superman lived on Krypton where the acceleration due to gravity is 23.6 m/s2. If a woman falls from a building that is 88.3 m tall, how long does he have to save her before she hits the ground? Im confused as to what equation I am to use. I know of the acceleration, initial velocity, as well the displacement. But what equation would I use?

563blackghost (563blackghost):

@Astrophysics

OpenStudy (astrophysics):

Do you know your kinematic formulas?

563blackghost (563blackghost):

Yes I do, but not by heart.

OpenStudy (astrophysics):

Alright so can you list them please :)

563blackghost (563blackghost):

Sure. \(\Large{d=v_{i}t +\frac{1}{2}at^{2}}\) \(\Large{v_{f}^{2}=v_{i}^{2}+2ad}\) \(\Large{v_{f}=v_{i}+at}\) \(\Large{d=\frac{v_{i}+v_{f}}{2}}t\)

OpenStudy (astrophysics):

Great ok, so what are we solving for here?

OpenStudy (stamp):

Superman lived on Krypton where the acceleration due to gravity is 23.6 m/s2. If a woman falls from a building that is 88.3 m tall, how long does he have to save her before she hits the ground? —————————————————————— a = 23.6 m/ss, Vi = 0, d = 88.3 m, Solve t.

563blackghost (563blackghost):

Well we are trying to find the time it takes before the woman hits the ground. We are well aware of the acceleration (23.6 m/s^2), the displacement (88.3m), and of the initial velocity (0 m/s).

OpenStudy (stamp):

Since Vi = 0, d = t(ta)/2, Solve t. 2d = t(ta) 2d/a = t(t) t = +/- sq(2d/a) Note: Since t is time we will accept only the positive value.

563blackghost (563blackghost):

So we would change the first equation into a function of t?

OpenStudy (stamp):

ab = c tells us what c is, and also tells us a = b/c or b = c/a. We use which depending on if we want a, b, or c. Our formula tells us distance d, but we rearrange it to tell us time t = +/- sq(2d/a)

563blackghost (563blackghost):

Ok. So my equation would be... \(\Large{t=(\frac{2(88.3)}{23.6})^{2}}\) right?

OpenStudy (stamp):

sq = square root, I understand the confusion. Let's call it t = root(2d/a)

563blackghost (563blackghost):

ooo ok i square root it... \(\Large{t=\sqrt{\frac{2(88.3}{23.6}}}\)

OpenStudy (stamp):

calculator time ... time t is expressed in the unit seconds s. So t = sqroot(2d/a) seconds Do you have the answer to cross-reference what you obtain on your calculator?

563blackghost (563blackghost):

I got 2.73 seconds. Would that be correct?

OpenStudy (stamp):

calculator doesn't lie

563blackghost (563blackghost):

Thank you so much! I was stuck on this for a while, but you got me out of it :D

OpenStudy (stamp):

don't sweat it and stay studying

563blackghost (563blackghost):

Will do!

OpenStudy (stamp):

If we did have an initial velocity Vi not equal to 0, our equation is d = tVi + at(t)/2 In which case, to solve for t, we would have to set the equation to 0 so that at(t)/2 + tVi - d = 0 Giving us a solvable quadratic equation t = quad. form. in which Qa = a, Qb = Vi, Qc = -d Quadratic Form of an Equation: Qa x^2 + Qb x + Qc = 0